### v1

@@ Acknowledgements
The authors are grateful to E.F. Grekova, S.D. Liazhkov, Yu.A. Mochalova, A.V. Porubov for useful and stimulating discussions.
Funding
The research is supported by the Ministry of Science and Higher Education of the Russian Federation (project 124040800009-8).
@@ Appendix A Non-dimensionalization
$m_{\infty}\frac{\mathrm{d}^{2}{\tilde{u}}_{n}}{\mathrm{d}\tilde{t}^{2}}-K_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})-(\tilde{u}_{n}-\tilde{u}_{n-1})\big{)}\\ -B_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})^{3}-(\tilde{u}_{n}-\tilde{u}_{n-1})^{3}\big{)}\\ =\Big{(}(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})-(\tilde{u}_{0}-\tilde{u}_{-1})\big{)}\\ +\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}-(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\big{)}\Big{)}\delta_{n}\\ -(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})\delta_{n-1}-(\tilde{u}_{0}-\tilde{u}_{-1})\delta_{n+1}\big{)}\\ -\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}\delta_{n-1}\\ -(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\delta_{n+1}\big{)}\\ -(\tilde{m}-m_{\infty})\frac{\mathrm{d}^{2}{\tilde{u}}_{0}}{\mathrm{d}\tilde{t}^{2}}\delta_{n}+\tilde{p}\delta_{n-N}\delta(\tilde{t}).$ (A.1)
Here, $n\in\mathbb{Z}$; $\tilde{t}$ is the time, $\tilde{u}_{n}(\tilde{t})$ is the displacement of the particle with a number $n$; $m_{\infty}$ is the mass for particles outside the defect; $K_{\infty}$, $B_{\infty}$ are the linear and cubic stiffnesses for the bonds outside the defect; $\tilde{m}$ is the mass of the particle at the defect, $\tilde{K}$ is the linear stiffness for the bonds at the defect; $\tilde{B}_{L}$ and $\tilde{B}_{R}$ are the cubic stiffnesses for the bonds to the left and to the right of the defect. Since the right-hand side of Eq. (A.1) involves the term proportional to the Dirac delta-function $\delta(t)$, we formulate initial conditions in the form, which is conventional in the theory of generalized functions (distributions):
$\tilde{u}_{n}\big{|}_{t<0}\equiv 0.$ (A.2)
@@
$\begin{gathered}\omega=\sqrt{\frac{K_{\infty}}{m_{\infty}}},\quad t=\omega\tilde{t},\quad u_{n}=\frac{\tilde{u}_{n}K_{\infty}}{\tilde{p}\omega},\\ K=\frac{\tilde{K}}{K_{\infty}},\quad m=\frac{\tilde{m}}{m_{\infty}},\\ B_{L}=\frac{\tilde{B}_{L}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad B_{R}=\frac{\tilde{B}_{R}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad\beta=\frac{B_{\infty}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}}.\end{gathered}$ (A.3)
Using dimensionless quantities (A.3), Eq. (A.1) may be rewritten in the following dimensionless form:
$\ddot{u}_{n}-(u_{n+1}-2u_{n}+u_{n-1})\\ -\beta\big{(}(u_{n+1}-u_{n})^{3}-(u_{n}-u_{n-1})^{3}\big{)}\\ =\big{(}(K-1)\big{(}(u_{1}-u_{0})-(u_{0}-u_{-1})\big{)}\\ +\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\big{)}\big{)}\delta_{n}\\ -(K-1)\big{(}(u_{1}-u_{0})\delta_{n-1}-(u_{0}-u_{-1})\delta_{n+1}\big{)}\\ -\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}\delta_{n-1}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\delta_{n+1}\big{)}\\ -(m-1)\ddot{u}_{0}\delta_{n}+\delta_{n-N}\,\delta(t),$ (A.4)
@@
which is equivalent to Eqs. (2.1)–(2.4) if notation (2.5) is accepted. The initial conditions (A.2) can be reformulated in the classical form (2.6).

### v2

@@ Acknowledgements
The authors are grateful to E.F. Grekova, S.D. Liazhkov, Yu.A. Mochalova, A.V. Porubov for useful and stimulating discussions. The research is supported by the Ministry of Science and Higher Education of the Russian Federation (project 124040800009-8).
@@ Appendix A Non-dimensionalization
$m_{\infty}\frac{d^{2}{\tilde{u}}_{n}}{d\tilde{t}^{2}}-K_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})-(\tilde{u}_{n}-\tilde{u}_{n-1})\big{)}\\ -B_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})^{3}-(\tilde{u}_{n}-\tilde{u}_{n-1})^{3}\big{)}\\ =\Big{(}(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})-(\tilde{u}_{0}-\tilde{u}_{-1})\big{)}\\ +\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}-(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\big{)}\Big{)}\delta_{n}\\ -(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})\delta_{n-1}-(\tilde{u}_{0}-\tilde{u}_{-1})\delta_{n+1}\big{)}\\ -\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}\delta_{n-1}\\ {-(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\delta_{n+1}\big{)}}\\ -(\tilde{m}-m_{\infty})\frac{d^{2}{\tilde{u}}_{0}}{d\tilde{t}^{2}}\delta_{n}.$ (20)
[…]
The initial conditions are
$\tilde{u}_{n}(0)=0,\quad\frac{d\tilde{u}_{n}}{d\tilde{t}}(0)=\frac{\tilde{p}\delta_{n-N}}{m_{\infty}+\delta_{n}(\tilde{m}-m_{\infty})}.$ (21)
Initial conditions in the form of Eq. (21) correspond to the pulse loading $\tilde{p}\delta(t)$ applied to the particle with label $N$. Here $\tilde{p}$ is the magnitude of the loading, $\delta(t)$ is the Dirac delta-function.
@@
$\begin{gathered}\omega=\sqrt{\frac{K_{\infty}}{m_{\infty}}},\quad t=\omega\tilde{t},\quad u_{n}=\frac{\tilde{u}_{n}K_{\infty}}{\tilde{p}\omega},\\ K=\frac{\tilde{K}}{K_{\infty}},\quad m=\frac{\tilde{m}}{m_{\infty}},\\ B_{L}=\frac{\tilde{B}_{L}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad B_{R}=\frac{\tilde{B}_{R}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad\beta=\frac{B_{\infty}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}}.\end{gathered}$ (22)
Using dimensionless quantities (22), Eq. (20) may be rewritten in the following dimensionless form:
$\ddot{u}_{n}-(u_{n+1}-2u_{n}+u_{n-1})\\ -\beta\big{(}(u_{n+1}-u_{n})^{3}-(u_{n}-u_{n-1})^{3}\big{)}\\ =\big{(}(K-1)\big{(}(u_{1}-u_{0})-(u_{0}-u_{-1})\big{)}\\ +\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\big{)}\big{)}\delta_{n}\\ -(K-1)\big{(}(u_{1}-u_{0})\delta_{n-1}-(u_{0}-u_{-1})\delta_{n+1}\big{)}\\ -\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}\delta_{n-1}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\delta_{n+1}\big{)}\\ -(m-1)\ddot{u}_{0}\delta_{n},$ (23)
@@
which is equivalent to Eqs. (1)–(4) if notation (5) is accepted. The initial conditions (21) can be reformulated in the form of Eq. (6).
Appendix B An infinite linear chain with an isotopic defect
For the reader’s convenience, we provide here some important formulas obtained in Shishkina and Gavrilov (2023a) that describes the solution of the problem considered in Sect. III, Eqs. (10)–(14) in the linear case $\beta=0$. The exact solution for the particle velocities in the integral form is an inverse Fourier transform:
${v}_{n}=-\frac{\mathrm{i}}{2\pi}\left(\int_{\mathbb{P}}+\int_{\mathbb{S}}\right)\Omega\mathcal{G}_{n}(\Omega)\mathrm{e}^{-\mathrm{i}\Omega t}\,d\Omega+\mathrm{c.c.}\\ =I_{n}^{\mathrm{pass}}+I_{n}^{\mathrm{stop}}+\mathrm{c.c.}=v_{n}^{\mathrm{pass}}+v_{n}^{\mathrm{stop}},$ (24)
$\mathcal{G}_{n}(\Omega)=-\frac{\mathrm{e}^{\mathrm{i}|n|\operatorname{sign}\Omega\,\arccos\frac{2-\Omega^{2}}{2}}}{(m-1)\Omega^{2}+\mathrm{i}\Omega\sqrt{4-\Omega^{2}}},\ \ \Omega\in\mathbb{P};$ (25)
$\mathcal{G}_{n}(\Omega)=\frac{(-1)^{|n|}2^{|n|}}{\Phi^{|n|-1}(\Omega)\big{(}(-m\Omega^{2}+2)\Phi(\Omega)+4\big{)}},\ \ \Omega\in\mathbb{S};$ (26)
$\mathrm{c.c.}$ are the complex conjugate terms. Here,
$\Phi(\Omega)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\Omega^{2}-2+|\Omega|\sqrt{\Omega^{2}-4};$ (27)
$\mathbb{P}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}[0,\Omega_{\ast}],\qquad\mathbb{S}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}(\Omega_{\ast},\infty)$ (28)
Remark 2.
In Shishkina and Gavrilov (2023a) mostly the heat, i.e., the kinetic energy, transfer processes are under consideration. Thus, the solution is obtained for the particle velocities, not for the displacements. This is also useful because the displacements involve a non-vanishing non-oscillating component (a rigid body mode of motion).
$0<m<1.$ (29)
If the last inequality is true, in the interval $(0,\Omega_{\ast})$ there exists a simple root
$\Omega_{0}=\frac{2}{\sqrt{m(2-m)}}$ (30)
of the denominator for $\mathcal{G}^{\mathrm{stop}}$. Therefore, integral $I^{\mathrm{stop}}$ does not exist in the classical sense. In the latter case, integral $I^{\mathrm{stop}}$ should be considered as the Fourier transform for a generalized function. To regularize this, we can apply Shishkina et al. (2023); Shishkina and Gavrilov (2023a) the limit absorption principle. Finally, for $t\to\infty$ at a fixed position $n$ one gets
$v_{n}^{\mathrm{stop}}=I^{\mathrm{stop}}+\mathrm{c.c.}=H(1-m)L_{n}(t)+O(t^{-1}),$ (31)
${L}_{n}=\frac{2(m-1)(-1)^{|n|+1}m^{|n|-1}H(1-m)}{(2-m)^{|n|+1}}\cos\Omega_{0}t,$ (32)
where $t\to\infty$, see Shishkina and Gavrilov (2023a). Here $L_{n}(t)$ is the localized non-vanishing oscillation, $H(\cdot)$ is the Heaviside step-function.
The phenomenon of the anti-localization is mostly related to the integral $I^{\mathrm{pass}}$, which describes the propagating part of the wave-field. Following to Gavrilov (2022), we have estimated it in Shishkina and Gavrilov (2023a) on a moving at an arbitrary sub-critical speed $w$ point of observation. Taking into account that solution (24) clearly is an even function of $n$, put
$|n|=wt,$ (33)
$0<w<1$ (34)
in the expression for $I^{\mathrm{pass}}$ defined by Eq. (24). The obtained integral can be estimated for $t\to\infty$ using the method of stationary phase. This provides an asymptotic continuum solution:
$v_{n}^{\mathrm{pass}}=I^{\mathrm{pass}}+\mathrm{c.c.}\\ =\frac{{A}(w)}{\sqrt{t}}\cos\Big{(}\phi(w)\,t+\frac{\pi}{4}+\psi(w)\Big{)}+O(t^{-3/2}),$ (35)
[…]
$\phi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}2(w\arccos w-\sqrt{1-w^{2}}),$ (40)
$\psi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\arctan\frac{(m-1)\sqrt{1-w^{2}}}{w}.$ (41)
Note that for $w>1$ there is no stationary point, and the corresponding term of order $1/\sqrt{t}$ is zero. This fact is taken into account by introducing the multiplier H(1w)H(1-w) in the numerator of the right-hand side of Eq. (39). In any system with an isotopic impurity, i.e., in the case (11), we have:
${A}(w)=\frac{\sqrt{2}w}{\sqrt{\pi}|m-1|}+O(w^{3}),\quad w\to 0;\qquad A(0)=0;$ (42)
and, thus, the amplitude $A(w)$ of the propagating part $v^{\mathrm{pass}}$ for the particle velocities is small in a certain expanding (since $w=|x|/t$) neighborhood of the impurity. We call this phenomenon the anti-localization of non-stationary waves. Finally, the asymptotic expansion for the right-hand side of Eq. (24) at $n=0$ (just at the impurity) has the following form
$v_{0}(t)=H(M-K)L_{0}(t)+\frac{\cos\left(2t-\frac{3\pi}{4}\right)}{2\sqrt{\pi}(m-1)^{2}\,t^{3/2}}\\ +o(t^{-3/2}).$ (43)
The first term in the right-hand side of Eq. (43) is the contribution from the poles $\pm\Omega_{0}$ and describes localized oscillation, which exists if and only if $m<1$. The second term is the anti-localized part of the wave-field at the inclusion expressed as the total contribution from the cut-off frequency $\Omega_{\ast}$ for both integrals $I^{\mathrm{stop}}$ and $I^{\mathrm{pass}}$.
$A(w)=\frac{1}{\sqrt{\pi{\sqrt{1-w^{2}}}}}=\frac{1}{\sqrt{\pi}}+O(w^{2}),\quad w\to 0;\\ \qquad A(0)\neq 0.$ (44)
Though the last formula can be formally derived from Eq. (39), this way is incorrect since Eq. (39) is obtained assuming Eq. (11). The correct way to obtain formula (44) is to use the exact solution (9) Gavrilov (2022). Thus, the anti-localization in the linear system is an asymptotic weakening of the propagating component of the wave-field. The order of decay for the vanishing component is $1/\sqrt{t}$ in the uniform system and $1/\sqrt{t^{3}}$ in the system with a defect (near the defect).
Appendix C A semi-infinite linear chain
The exact solution is Liazhkov (2025)
$v_{n}=J_{2(n-1)}(2t)+J_{2n}(2t),\qquad n\in\mathbb{N}.$ (45)
The anti-localization in the semi-infinite linear chain ($\beta=0$) was demonstrated in study Gavrilov et al. (2023), where the asymptotic solution for $v_{n}$ in the form of the right-hand side of Eq. (35) wherein
${A}(w)=\frac{2w}{\sqrt{\pi}\sqrt[4]{1-w^{2}}},$ (46)
$\psi=\arctan\frac{\sqrt{1-w^{2}}}{w},$ (47)
$w=\frac{|n-1|}{t}$ (48)
was obtained. The expression for $\phi$ coincides with (40). Clearly, $A(0)=0$, which implies the anti-localization effect. For the end particle, one has the anti-localized solution Gavrilov et al. (2023):
$v_{1}=\frac{\sin(2t-\frac{\pi}{4})}{\sqrt{\pi}t^{3/2}}+o(t^{-3/2}).$ (49)